# New Foundations is consistent

### v22

@@ 4.4 Impredicativity: verifying the axiom of union
Suppose that $\{\{g\}_{\beta}:g\in G\}_{\alpha}$ is symmetric. It then has a strong support $S$. We claim that $S_{(\beta)}$, defined as $\{(z,C):{\tt max}(C)=\beta\wedge(z,C\cup\{\alpha\})\in S\}$, is a $\beta$-support for $G_{\beta}$.
@@
while at the same time $(\pi^{*}_{\beta})_{\gamma}$ agrees with $\pi_{\gamma}$ on $G$. This implies that $\pi_{\gamma}(g)\in G$ (and the same argument applies to $\pi^{-1}$) so $\pi$ fixes $G_{\beta}$’.
@@
We construct the allowable permutation $\pi^{*}$ by the Freedom of Action Theorem (4.18), by defining the following $\alpha$-approximation $\psi$.
@@
For every ($(\{x\},A)\in{\tt rng}(S)$, we define $\psi_{A}(x)=x$, and for every $(N,A)\in{\tt rng}(S)$ for $N$ a near-litter, we define $\psi_{A}(N^{\circ})=N^{\circ}$. Extend $T^{\uparrow\beta}$ its canonical strong $\beta$-support $T^{*}$ (remark 4.23), then if ($(\{x\},A)\in{\tt rng}(T^{*})$, we define $(\psi_{\beta})_{A}(\pi^{n}_{A}(x))=\pi^{n+1}_{A}(x)$ for every integer $n$, and if $(N,A)\in{\tt rng}(T^{*})$ for $N$ a near-litter, we similarly define $(\psi_{\beta})_{A}(\pi^{n}_{A}(N)^{\circ})=\pi^{n+1}_{A}(N)^{\circ}$.
We then complete orbits of atoms similarly to the proof of lemma 4.29 to ensure that $\psi_{A}(N)=N$ for $(N,A)\in{\tt rng}(S)$ and $(\psi_{\beta})_{A}(\pi^{n}_{A}(N))=\pi^{n+1}_{A}(N)$ for $(N,A)\in{\tt rng}(T^{*})$, and then arrange extra orbits if needed such that the interference condition in definition 4.1 is satisfied.
@@
It now suffices to check that this approximation is coherent. If $\psi_{A}^{*}``L\sim L^{\prime}$, we have two cases: either $(N,A)$ appears in $S$ with $N^{\circ}=L=L^{\prime}$, or $B=A\setminus\{\alpha\}$ has maximum element $\beta$ and $(N,B)$ occurs in $T^{*}$ with $\pi_{B}^{n}(N)^{\circ}=L$ and $\pi_{B}^{n+1}(N)^{\circ}=L^{\prime}$. Note that if both of these cases are true at the same time, then as $\pi$ fixes $S_{(\beta)}$, the two defined images of $L$ coincide.
@@
In the first case, if $L$ is $A$-flexible then $L^{\prime}=L$ is $A$-flexible as required, and if $L=f_{\delta,{\tt min}(A_{1})}(x,T)$ with $\delta<{\tt min}(A_{2})$, then as $S$ is a strong support, the range of $T^{\uparrow A_{2}\cup\{\delta\}}$ is a subset of the range of $S$, as required.
@@
For the second case, we first show that $L$ is $A$-flexible if and only if $L$ is $B$-flexible. The forward direction is trivial. The converse can only fail if $B$ is too short to satisfy the definition of inflexibility, so we may suppose that $L=f_{\delta,{\tt min}(A_{1})}(x,T)$ with $\delta<{\tt min}(A_{2})$ where $A$ has length exactly 3. Then we must have ${\tt min}(A_{1})=\beta$, so $A=\{\alpha,\beta,-1\}$ and $B=\{\beta,-1\}$. As $(N,B)$ occurs in $T^{*}$, it either came from a support condition in $T$, which is impossible as $B$ does not contain $\gamma$, or it came from a type-raised copy of an $\eta$-support we needed to include in $T^{*}$ to make it a strong support. But $\eta\neq-1$ as $-1$-supports cannot contain near-litter support conditions, and if $\eta$ were a proper type index, we would have $-1<\eta<\beta$ and $\eta\in B$, also giving a contradiction. Hence $(T^{*})^{\uparrow\alpha}$ is a strong $\alpha$-support, and we are done by the method used for $S$.
@@
This completes the proof. In the formal proof in Lean, what is actually done is a proof that each of the assertions in the finite axiomatization of Hailperin in the version discussed in subsection 2.5 holds in all typed versions in our structure for the language of TTT, so it is in fact a model of TTT. The axioms in Hailperin other than the axiom of type lowering are predicative comprehension axioms and admit demonstration by the methods of proposition 4.34 [or] section 4.3, done explicitly without metamathematics. In the formalization, the axiom of type lowering, which contains rather more content than the axiom of set union restricted to sets of singletons which is proved here, is proved by first proving the existence of an iterated image under elementwise application of the singleton operation of the desired set, whose definition is predicative, then repeatedly applying the result of this section that sets of singletons have unions.
@@ 5 Conclusions, extended results, and questions
This is a rather boring resolution of the NF consistency problem.
@@
Questions of consistency with NF of global choice-like statements such as “the universe is linearly ordered” cannot be resolved by the method used here (at least, not without major changes). One statement which seems to be about big sets can be seen to hold in our models: the power set of any well-orderable set is well-orderable, and more generally, beth numbers are alephs. We indicate the proofs: a relation which one of our models of TTT thinks is a well-ordering actually is a well-ordering, because the models are countably complete; so a well-ordering with a certain support has all elements of its domain sets with the same support (a permutation whose action fixes a well-ordering has action fixing all elements of its domain), and all subsets of and relations on the domain are sets with the same support (adjusted for type differential), and this applies further to the well-ordering of the subsets of the domain which we find in the metatheory. Applying the same result to sets with well-founded extensional relations on them proves the more general result about beth numbers. This form of choice seems to allow us to use choice freely on any structure one is likely to talk about in the usual set theory. It also proves, for example, that the power set of the set of ordinals (a big set!) is well-ordered.
@@
NF with strong axioms such as the Axiom of Counting (introduced by Rosser in [14], an admirable textbook based on NF), the Axiom of Cantorian Sets (introduced in [5]) [^8] Getting Cantorian Sets or Large Ordinals to hold is very sensitive to the relationship between $\kappa$ and $\lambda$, and the first author is not yet entirely certain of the details. It requires the hypothesis that $\kappa<\lambda=\mu$ to avoid outright refutability of this axiom in resulting models of NF, and then a large cardinal hypothesis. We note that for the moment we need to use Henson’s original formulation of the Axiom of Cantorian Sets, that any well-orderable cantorian set is strongly cantorian; the situation for non-well-orderable sets is less clear. or my axioms of Small Ordinals and Large Ordinals (introduced in my [7], which pretends to be a set theory textbook based on NFU) can be obtained by choosing $\lambda$ large enough to have strong partition properties, more or less exactly as I report in my paper [8] on strong axioms of infinity in NFU: the results in that paper are not all mine, and I owe a good deal to Robert Solovay in that connection (unpublished conversations and [18]).
@@
That NF has $\alpha$-models for each standard ordinal $\alpha$ should follow by the same methods Jensen used for NFU in his original paper [10]. No model of NF can contain all countable subsets of its domain; all well-typed combinatorial consequences of closure of a model of TST under taking subsets of size $<\kappa$ will hold in our models, but the application of compactness which gets us from TST + Ambiguity to NF forces the existence of externally countable proper classes, a result which has long been known and which also holds in NFU.
@@
We mention some esoteric problems which our approach solves. The Theory of Negative Types of Hao Wang (TST with all integers as types, proposed in [22]) has $\omega$-models; an $\omega$-model of NF gives an $\omega$-model of the theory of negative types immediately. The question of existence of $\omega$ models of the theory of negative types was open.
@@
In ordinary set theory, the Specker tree of a cardinal is the tree in which the top is the given cardinal, the children of the top node are the preimages of the top under the map $(\kappa\mapsto 2^{\kappa})$, and the part of the tree below each child is the Specker tree of the child. Forster proved using a result of Sierpinski that the Specker tree of a cardinal must be well-founded (a result which applies in ordinary set theory or in NF(U), with some finesse in the definition of the exponential map in NF(U)). Given Choice, there is a finite bound on the lengths of the branches in any given Specker tree. Of course by the Sierpinski result a Specker tree can be assigned an ordinal rank. The question which was open was whether existence of a Specker tree of infinite rank is consistent. It is known that in NF with the Axiom of Counting the Specker tree of the cardinality of the universe is of infinite rank. Our results in this paper can be used to show that Specker trees of infinite rank are consistent in bounded Zermelo set theory with atoms or without foundation (this takes a little work, using the way that internal type representations unfold in TTT and a natural interpretation of bounded Zermelo set theory in TST; a tangled web as described above would have range part of a Specker tree of infinite rank). A bit more work definitely gets this result in ZFA, and we are reasonably confident that our permutation methods can be adapted to ZFC using forcing in standard ways (in which we are not expert) to show that Specker trees of infinite rank can exist in ZF.

### v23

@@ 4.6 Impredicativity: verifying the axiom of union
Suppose that $\{\{g\}_{\beta}:g\in G\}_{\alpha}$ is symmetric. It then has a strong support $S$. We claim that $S_{(\beta)}$, defined as $\{(z,C):{\tt max}(C)=\beta\wedge(z,C\cup\{\alpha\})\in S\}$, is a $\beta$-support for $G_{\beta}$, where without loss of generality we assume that $S$ has been reordered such that the domain of $S_{(\beta)}$ (and $(S_{(\beta)})_{(\gamma)}$) is an ordinal.
@@
while at the same time $(\pi^{*}_{\beta})_{\gamma}$ agrees with $\pi_{\gamma}$ at $g$. This implies that
$\pi_{\gamma}(g)=(\pi^{*}_{\beta})_{\gamma}(g)\in(\pi^{*}_{\beta})_{\gamma}``G=G.$
The same construction can be done for $\pi^{-1}$ in place of $\pi$ (for each $g$), so $\pi$ fixes $G_{\beta}$.
@@
We construct the allowable permutation $\pi^{*}$ by the Freedom of Action Theorem (4.23), by defining the following $\alpha$-action $\xi$.
@@
For every ($(\{x\},A)\in{\tt rng}(S)$, we define $\xi_{A}(x)=x$, and for every $(N,A)\in{\tt rng}(S)$ for $N$ a near-litter, we define $\xi_{A}(N)=N$. Extend $T^{\uparrow\beta}$ to a strong $\beta$-support $T^{*}$ (as in proposition 4.26), then if ($(\{x\},B)\in{\tt rng}(T^{*})$, we define $(\xi_{\beta})_{B}(x)=\pi_{B}(x)$, and if $(N,B)\in{\tt rng}(T^{*})$ for $N$ a near-litter, we similarly define $(\xi_{\beta})_{B}(N)=\pi_{B}(N)$.
It now suffices to check that this action is coherent; then, by the Freedom of Action Theorem, we will obtain $\pi^{*}$ with the property that $\pi^{*}[S]=S$ and $\pi^{*}_{\beta}[T^{*}]=\pi[T^{*}]$, so in particular, $(\pi^{*}_{\beta})_{\gamma}[T]=\pi_{\gamma}[T]$.
@@
If $\xi_{A}(N)$ is defined, we have two cases: either $(N,A)$ appears in $S$, in which case $\xi_{A}(N)=N$, or $B=A\setminus\{\alpha\}$ has maximum element $\beta$ and $(N,B)$ occurs in $T^{*}$ with $\xi_{A}(N)=\pi_{B}(N)$. Note that if both of these cases are true at the same time, then as $\pi$ fixes $S_{(\beta)}$, the two defined images of $N$ coincide.
@@
In the first case, the case for flexibility is trivial, and if $N\sim f_{\delta,{\tt min}(A_{1})}(x,U)$ with $\delta<{\tt min}(A_{2})$, then as $S$ is a strong support, the range of $U^{\uparrow A_{2}}$ is a subset of the range of $S$, as required.
We now consider the second case. If $N$ is $A$-flexible, we claim that $\pi_{B}(N)$ is also $A$-flexible. Suppose that $\pi_{B}(N)\sim f_{\delta,\gamma}(x,U)$ for $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$. Then by the coherence condition, assuming $B$ has length at least 3,
$N\sim\pi_{B}^{-1}``f_{\delta,\gamma}(x,U)\sim f_{\delta,\gamma}((\pi_{B_{2}}^{-1})_{\delta}(x),(\pi_{B_{2}}^{-1})_{\delta}[U]),$
@@
so $N$ is $A$-inflexible, giving a contradiction. In the case that $B$ does not have length at least 3, it must have length exactly 2 because $A$ has length at least 3, so we have $A=\{\alpha,\beta,-1\}$ and $B=\{\beta,-1\}$. Note that $(N,B)$ occurs in $T^{*}$, so it either came from a support condition in $T$, which is impossible as $B=\{\beta,-1\}$ does not contain $\gamma$, or it came from a type-raised copy of an $\eta$-support we needed to include in $T^{*}$ to make it a strong support. But $\eta\neq-1$ as $-1$-supports cannot contain near-litter support conditions, and if $\eta$ were a proper type index, we would have $-1<\eta<\beta$ and $\eta\in B$, also giving a contradiction.
Finally, suppose that $N$ is $A$-inflexible, so $N\sim f_{\delta,\gamma}(x,U)$ for $\gamma={\tt min}(A_{1})$ and $\delta<{\tt min}(A_{2})$. By the argument of the previous paragraph, we must conclude that $A$ is of length at least 4, and so $\gamma={\tt min}(B_{1})$ and $\delta<{\tt min}(B_{2})$. As $T^{*}$ is strong, it contains a suitably type-raised copy of $U$, and so we are done by definition of the action of $\xi$ on $T^{*}$.
@@
This completes the proof. In the formal proof in Lean, what is actually done is a proof that each of the assertions in the finite axiomatization of Hailperin in the version discussed in subsection 2.4 holds in all typed versions in our structure for the language of TTT, so it is in fact a model of TTT. The axioms in Hailperin other than the axiom of type lowering are predicative comprehension axioms and admit demonstration by the methods of section 4.5, done explicitly without metamathematics. In the formalization, the axiom of type lowering, which contains rather more content than the axiom of set union restricted to sets of singletons which is proved here, is proved by first proving the existence of an iterated image under elementwise application of the singleton operation of the desired set, whose definition is predicative, then repeatedly applying the result of this section that sets of singletons have unions.
@@ 5 Conclusions, extended results, and questions
In a series of subsections, we discuss further results supported by our model construction and further questions.
5.1 Ways in which this resolution of the NF consistency problem is unexciting
This is in some ways a rather boring resolution of the NF consistency problem.
@@ 5.1 Ways in which this resolution of the NF consistency problem is unexciting
5.2 Questions and a result about global choice-like statements
Questions of consistency with NF of global choice-like statements such as “the universe is linearly ordered” or the prime ideal theorem cannot be resolved by the method used here (at least, not without major changes). Our models falsify the existence of a global linear order at the outset: a litter cannot be linearly ordered. We see no reason to believe that the universe cannot be linearly ordered in NF or that the prime ideal theorem cannot hold, but we do not know how to model these situations.
We do have one quite interesting theorem which appears to show that most practical applications of the axiom of choice are supported in our models of TTT and so with care in the resulting models of NF.
Note that in our models of TTT, any relation which the internal theory thinks is a well-ordering actually is a well-ordering. The reason is that in the metatheory any linear order which is not a well-ordering admits a countable set with no minimal element (the range of a strictly decreasing sequence). This requires choice, but choice holds in the metatheory. So, if a relation represented in the model of TTT is not a well-ordering from the standpoint of the metatheory, there is a countable subcollection of its domain which has no minimal element, and this countable subset is actually realized in the model of TTT, because our models of TTT are countably complete, so the model of TTT sees that the relation is not a well-ordering.
Essentially the same argument shows that any relation which a model of TTT constructed by our method thinks is well-founded is actually well-founded in the sense of the metatheory.
Theorem:
In our models of TTT, the power set of any well-orderable set can be well-ordered.
Proof:
Suppose that $X$ is a set in a model of TTT constructed by our method which admits a well-ordering (defined along a particular type sequence). The well-ordering $\leq_{X}$ of $X$ which we are given must have a support. A permutation which fixes the support of $\leq_{X}$ [in the appropriate sense defined in the paper] will fix $\leq_{X}$ and, because $\leq_{X}$ is a well-ordering in the sense of the metatheory, its action on the type of the elements of $X$ (in the type sequence with respect to which $\leq_{X}$ is a well-ordering) must fix each element of $X$, so in fact we have a common support for all elements of $X$, which in turn gives a support for each subcollection of $X$ in the sense of the metatheory as implemented in the appropriate type in the particular type sequence, so all of these are sets in the model of TTT (in the appropriate type in the particular type sequence) and further gives a support for the well-ordering of the power set of $X$ induced by any well-ordering of the power set of $X$ in the sense of the metatheory, so the power set of $X$ is well-ordered in the model of TTT.
It is worth remarking that this theorem has been proved in the Lean formalization.
It is an interesting incidental remark that the theorem is equivalent in NF to the assertion that a single big power set, the power set of the set of ordinals, can be well-ordered, and this is seen to hold in models of NF constructed from our models of TTT.
This form of choice seems to support all practical applications of choice in mathematics outside of set theory, and a lot of the applications inside set theory.
Applying the same technique to sets with well-founded extensional relations on them can be used to show that all $\beth$ numbers are alephs, and that the internal representation of an initial segment of the cumulative hierarchy using equivalence classes of well-founded extensional relations with top which is possible in TTT will agree precisely with the cumulative hierarchy in the metatheory as far as it goes (the internal representation will always give a proper initial segment of the hierarchy in the metatheory, of course).
5.3 Strong unstratified axioms for NF derived from strong partition properties
@@
NF with strong axioms such as the Axiom of Counting (introduced by Rosser in [14], an admirable textbook based on NF), the Axiom of Cantorian Sets (introduced in [5]) [^13] Getting Cantorian Sets or Large Ordinals to hold is sensitive to the relationship between $\kappa$ and $\lambda$. It requires the hypothesis that $\kappa<\lambda=\mu$ to avoid outright refutability of this axiom in resulting models of NF, and then a large cardinal hypothesis. We note that for the moment we need to use Henson’s original formulation of the Axiom of Cantorian Sets, that any well-orderable cantorian set is strongly cantorian; the situation for non-well-orderable sets is not yet clear to us. or my [^14] The first author is speaking. axioms of Small Ordinals and Large Ordinals (introduced in my [7], which pretends to be a set theory textbook based on NFU) can be obtained by choosing $\lambda$ large enough to have strong partition properties, more or less exactly as I report in my paper [8] on strong axioms of infinity in NFU: the results in that paper are not all mine, and I owe a good deal to Robert Solovay in that connection (unpublished conversations and [18]).
That NF has $\alpha$-models for each standard ordinal $\alpha$ follows by the same methods Jensen used for NFU in his original paper [10]: the proof we give in subsection 2.3.1 for the existence of $\alpha$-models is essentially the same as Jensen’s, and a model of TTT constructed by our method will implement each ordinal $\alpha$ of the metatheory which is less than $\kappa$ (and we can choose $\kappa$ larger than any fixed $\alpha$ and $\lambda$ with cofinality greater than $|\alpha|^{|\alpha|}$).
@@
No model of NF can contain all countable subsets of its domain; all well-typed combinatorial consequences of closure of a model of TST under taking subsets of size $<\kappa$ will hold in our models, but the application of compactness which gets us from TST + Ambiguity to NF forces the existence of externally countable proper classes, a result which has long been known and which also holds in NFU.
5.4 Some esoteric questions answered
@@
We mention some esoteric problems which our approach solves. The Theory of Negative Types of Hao Wang (TST with all integers as types, proposed in [22]) has $\omega$-models; an $\omega$-model of NF gives an $\omega$-model of the theory of negative types immediately. The question of existence of $\omega$-models of the theory of negative types was open.
@@
In ordinary set theory, the Specker tree of a cardinal is the tree in which the top is the given cardinal, the children of the top node are the preimages of the top under the map $(\kappa\mapsto 2^{\kappa})$, and the part of the tree below each child is the Specker tree of the child. Forster proved using a result of Sierpiński that the Specker tree of a cardinal must be well-founded (a result which applies in ordinary set theory or in NF(U), with some finesse in the definition of the exponential map in NF(U)). Given Choice, there is a finite bound on the lengths of the branches in any given Specker tree. Of course by the Sierpiński result a Specker tree can be assigned an ordinal rank. The question which was open was whether existence of a Specker tree of infinite rank is consistent. It is known that in NF with the Axiom of Counting the Specker tree of the cardinality of the universe is of infinite rank. Our methods show the existence of a model of NF with the Axiom of Counting. This gives us a model of TST in which the Specker tree of some cardinal is of infinite rank. It is straightforward to produce a model of bounded Zermelo set theory either with atoms or with extensionality and failure of foundation from a model of TST, which will satisfy the same stratified sentences: identify the elements of type 0 either as atoms or with their own singletons, then implement elements of each type $i+1$ as subsets of type $i$ as guided by the original model of TST, identifying type $i+1$ objects with elements of lower types which have been assigned the same extension. This shows that we can get a model of bounded Zermelo set theory with atoms or without foundation in which there is a cardinal with Specker tree of infinite rank. It is straightforward with our methods to get a model of TTT with the Axiom of Counting with inaccessibles cofinal in $\lambda$, and in this model we will have an inaccessible greater than a cardinal with Specker tree of infinite rank, and so a model of ZFA in which there is a cardinal with Specker tree of infinite rank. It’s actually a bit easier than this: the existence of a cardinal with Specker tree of infinite rank holds in any model built by our method in which $\lambda>\omega$, but referring back to the result about NF + Counting makes it easier to give a brief account here.
The further question remains as to whether it is consistent with ZF that there is a cardinal with Specker tree of infinite rank. We would guess that this could be shown to be true using symmetric forcing extensions, but we are not expert at this and leave the question to others.
5.5 The question of consistency strength
@@
5.6 The wide open question: what do models of NF look like in general?